# Zero-shot generalization across architectures for visual classification

### v2

@@ Abstract
Generalization to unseen data is a key desideratum for deep networks, but its relation to classification accuracy is unclear. Using a minimalist vision dataset and a measure of generalizability, we show that popular networks, from deep convolutional networks (CNNs) to transformers, vary in their power to extrapolate to unseen classes both across layers and across architectures. Accuracy is not a good predictor of generalizability, and generalization varies non-monotonically with layer depth. Code is available at github.com/dyballa/zero-shot-generalization.
@@ A.2 CIFAR-100 dataset
Figure 4: Generalization across fine-tuning epochs for all studied networks. We plot $g_{\mathrm{seen}}$ and test-set accuracy (using seen classes) for reference. Notice how $g_{\mathrm{seen}}$ is higher for all networks, as expected. This demonstrates that our $g$ metric does capture the ability of the networks’ to generalize to unseen classes.
Figure 5: The generalizability trend across layers may vary considerably from model to model. Although deeper layers tend to do better separate the examples from unseen classes, $g_{i}$ does not increase monotonically with depth, and for most models (ResNet, ViT, PViT, and PoolFormer) the best generalization is found at intermediary layers. This hold true for both datasets used for fine-tuning (calligraphy or CIFAR), showing that this phenomenon does not seem dataset-specific. Moreover, similar trends were found regardless of the metric used (NMI or $k$-nearest neighbors) as demonstrated by qualitatively similar curve shapes, despite being based on completely different methods. In almost all cases, the two metrics identified the same layer as the one that best generalized to the unseen classes. Interestingly, the $g_{i}$ curves are also qualitatively similar across datasets, indicating that the patterns observed are due to the network’s architecture, and not the dataset.

### v3

@@ Abstract
Generalization to unseen data is a key desideratum for deep networks, but its relation to classification accuracy is unclear. Using a minimalist vision dataset and a measure of generalizability, we show that popular networks, from deep convolutional networks (CNNs) to transformers, vary in their power to extrapolate to unseen classes both across layers and across architectures. Accuracy is not a good predictor of generalizability, and generalization varies non-monotonically with layer depth. Our code is available at here.
@@ A.2 CIFAR-100 dataset
Figure 4: Generalization across fine-tuning epochs for all studied networks. We plot $g_{\mathrm{seen}}$ and test-set accuracy (using seen classes) for reference. Notice how $g_{\mathrm{seen}}$ is higher for all networks, as expected. This demonstrates that our $g$ metric is able to capture the ability of the networks to generalize to unseen classes.
Figure 5: The generalizability trend across layers may vary considerably from model to model. Although deeper layers tend to better separate the examples from unseen classes, $g_{i}$ does not increase monotonically with depth, and for most models (ResNet, ViT, PViT, and PoolFormer) the best generalization is found at intermediary layers. This hold true for both datasets used for fine-tuning (calligraphy or CIFAR), showing that this phenomenon does not seem dataset-specific. Moreover, similar trends were found regardless of the metric used (NMI or $k$-nearest neighbors) as demonstrated by qualitatively similar curve shapes, despite being based on completely different methods. In almost all cases, the two metrics identified the same layer as the one that best generalized to the unseen classes. Interestingly, the $g_{i}$ curves are also qualitatively similar across datasets, indicating that the patterns observed are due to the network’s architecture, and not the dataset.